QUESTION FOUR “That old equipment for producing subassemblies is worn out”, said Mr. Paul Adom, president of Ebeyeyie Ghana Limited (EGL). “We need to make a decision quickly”. The company is trying to decide whether it should purchase new equipment and continue to make its subassemblies internally or whether it should discontinue production of its subassemblies and purchase them from an outside supplier. The alternatives are: Alternative 1: New equipment for producing the subassemblies can be purchased at a cost of GH₵1,750,000. The equipment would have a five-year useful life (the company uses straight-line depreciation method) and a GH₵250,000 salvage value. Alternative 2: The subassemblies can be purchased from an outside supplier who has offered to provide them for GH₵40 each under a five-year contract. EGL’s present costs per unit of producing the subassemblies internally (with the old equipment) are given below. These costs are based on a current activity level of 200,000 subassemblies per year: GH₵ Direct materials 13.75 Direct labour 20.00 Variable overhead 3.00 Fixed overhead (GH₵3.75 supervision, GH₵4.50 Depreciation, and GH₵10 general company overhead) 18.25 Total cost per unit 55.00 The new equipment would be more efficient and, according to the manufacturer, would reduce direct labour costs and variable overhead costs by 25%. Supervision cost (GH₵150,000 per year) and direct materials cost per unit would not be affected by the new equipment. The new equipment’s capacity would be 300,000 subassemblies per year. The company has no other use for the space now being used to produce subassemblies. REQUIRED: (A) The president is unsure what the company should do and would like an analysis showing what unit costs and what total costs would be under each of the two alternatives given above. Assume that 40,000 subassemblies needed each year. Which course of action would you recommend to the president and why? (B) Would your recommendation in (a) above be the same if the company’s needs were: (i) 250,000 subassemblies per year, or (ii) 300,000 subassemblies per year? Show computations in good form. (C) What qualitative factors would you recommend that the company consider before deciding?
Answer #1

DEPRECIATION OF NEW EQUIPMENT

PER YEAR = (1,750,000-250,000)/5=300,000

PER UNIT

IF 40,000= 300,000/40,000=7.5

IF 250,000= 300,000/250,000=1.2

IF 300,000= 1

IF NEW EQUPIT PURCHASE  would reduce direct labour costs and variable overhead costs by 25%

NEW COST OF DIRECT LABOUR =20 x 80% = 16/u

new variable oh cost = 3x80%= 2.4/u

supervision cost total 150,000

if 40,000 unit = 150,000/40000=3.75

250,000 = 150,000/250,000=.6

300,000 = 150,000/300,000=.5

THIS ALL IS BASED ON THE ASSUMPTION THAT SUPERVIOSRS SALARY ARE FIXED

A)

IF BUY MACHINE

NEW COST CARD

DIRECT MATERIAL =13.75

DIRECT LABOUR = 16

V OH = 2.4

FIXED COST= (3.75+7.5+10)=21.25

COST PER UNIT=53.4(13.5+16+2.4+21.25)

IF BUY FROM OUTSIDE

CAN BUY AT 40/ UNIT

I WILL RECOMEND TO BUY FROM OUT SIDE , COST PER UNIT OF BUYING LESS THAN PRODUCING INHOUSE(GIVE A SAVING PER UNIT OF 13.4/UNIT)

B)

(i) 250,000/YEAR

COST CARD

DIRECT MATERIAL =13.75

DIRECT LABOUR = 16

V OH = 2.4

FIXED COST= .6+1.2+10=11.8

COST PER UNIT=43.95

STILL WILL RECOMEND TO BUY, BUY PRICE IS 40

(ii) 300,000

COST CARD

DIRECT MATERIAL =13.75

DIRECT LABOUR = 16

V OH = 2.4

FIXED COST= .5+1+10=11.5

COST PER UNIT=43.65

STILL RECOMEND TO BUY

C)

Quality could arguably be the most important factor to consider when deciding to outsource. Quality can be hard enough to control in-house; however, when operations are outsourced, quality control becomes exponentially harder. It’s critical that you assess a potential manufacturing partner’s quality standards and understand how they’ll observe, track, and report on quality.

Quality is more than just a manufacturer’s certifications and standards, and can be observed by doing things like taking plant tours, looking at other projects the manufacturer is working on, and understanding their organizational culture. Quality standards and expectations are different from company to company and each product they produce, so when outsourcing it’s important that there’s agreement on all quality discussions.

While outsourcing demands vigilance and observation, it does provide the unique opportunity for manufacturers to upgrade their quality. For instance, if a manufacturer can’t fabricate a product to the necessary tolerances or if there’s not enough technical knowledge to perform specific tasks, then it could be possible to communicate to the potential manufacturer how making changes will improve their quality standards and would benefit them in the long term and open the door to future business. Quality control is still difficult as the supply chain becomes more complex, but with proper attention and diligence, there are opportunities to improve overall product quality.